2(4x + 15) = 8x + 3
Answer by funmath(2933)

2(4x + 15) = 8x + 3
2(4x)+2(15)=8x+3
8x+30=8x+3
-8x+8x+30=-8x+8x+3
30=3
The variables were eliminated and 30 does not equal 3. There is no solution.
Happy Calculating!!!